OpenStudy (anonymous):

need help solving (7)/(b-5) - (8)/(5-b)

OpenStudy (anonymous):

Look for the lowest common denominator, which is \[(b-5)*(5-b)\] Remember that you want the fraction to stay the same even after you multiply to get the lowest common denominator, so multiply by 1: \[\frac{5-b}{5-b} = \frac{b-5}{b-5} = 1\] \[(\frac{5-b}{5-b}*\frac{7}{b-5}) - (\frac{b-5}{b-5}*\frac{8}{5-b})\] \[\frac{(35-7b)-(8b-40)}{-b^{2}+10b-25}\] You can do the rest :)